A 4-pole shunt motor is lap-wound with 728 conductors, and has a flux of 25 mWb per

pole. The armature takes 50 Amp in current, the resultant demagnetizing effect due to

armature reaction reduces the air gap flux by 5%. What is developed torque of this

motor?


1
Expert's answer
2021-05-25T02:41:02-0400

"Torque=\frac{1}{2 \pi}\times\frac{ZP}{A}\times \phi I_a"

"Torque=\frac{1}{2 \pi}\times\frac{Z}{1}\times \phi I_a"

"Torque=\frac{1}{2 \pi}\times\frac{728}{1}\times 25 \times 10^{-3} \times 50"

"Torque=\frac{910}{2 \pi}"

Considering the resultant demagnetizing effect due to armature reaction reduces the air gap flux by 5%

"Torque=\frac{910}{2 \pi} \times \frac{95}{100}=137.6 N-m"